OpenStudy (anonymous):

1. If a function f(x) is shifted to the left one unit, what function represents the transformation? A.f(x+1) <---- i think its this one B.f(x-1) C.f(x)+1 D.f(x)-1 2.Let g(x) be the reflection of f(x)=x^2+5 in the x-axis. what is a function rule for g(x)? A.g(x)=-x^2-5 B.g(x)=x^2+5 C.g(x)=x^2-5 D.g(x)=-x^2+5<--- i think this is it 3. What transformations change the graph of f(x) to the graph of g(x)? f(x)=-7x^2 g(x)=-35x^2+5 4. write an equation for the following transformation of y=x: a vertical stretch by a factor of 6. A.y=6x B.y=x/6<---- i think C.y=x+6 D.y=x-6

OpenStudy (anonymous):

can someone help me with these i think i have the right answers but im not sure

OpenStudy (johnweldon1993):

1 is correct 2, if we reflect a function f(x) completely over the x-axis...it becomes -f(x) so if your function is f(x) = x^2 + 5....that would give us -f(x)=-x^2 - 5 3) where is your work? :D 4) To vertically stretch ...we multiply a function by a constant so f(x) would be "constant" times f(x) To vertically stretch a function, that constant needs to be > 1...so to vertically stretch a function f(x) = x by 6...we have y = 6x

OpenStudy (anonymous):

okay for number three my bad i thought i got it all in there give me a sec to put that on here.

OpenStudy (anonymous):

A. the graph of g(x) is the graph of f(x) stretched vertically by a factor of 1/5 and translated down by 5 units. B. the graph of g(x) is the graph of f(x) stretched vertically by a factor of 5 and translated down by 5 units C.the graph of g(x) is the graph of f(x) stretched vertically by a factor of 1/5 and translated up by 5 units. D.the graph of g(x) is the graph of f(x) stretched vertically by a factor of 5 and translated up by 5 units for this one i got b but im not to sure

OpenStudy (anonymous):

can someone help me with number 3????